OpenStudy (anonymous):

Can someone please help me on this :c I really need to understand this lesson before I take the test next week !

OpenStudy (anonymous):

PLEASEEEEEE, I attatched it down below.

OpenStudy (jdoe0001):

have you covered geometric sequences yet?

OpenStudy (anonymous):

@jdoe0001 sort of but I dont understand what the assignment wants me to do when answering the three questions at the bottom. Im not very good at worded out mathematical questions :c

OpenStudy (anonymous):

@zepdrix @Luigi0210 ?? can you help :c

zepdrix (zepdrix):

Hmm this problem doesn't look very fun ;c

OpenStudy (anonymous):

It isnt. I dont know where to start with it >_<

zepdrix (zepdrix):

So what is the common ratio, \(\Large\rm r\), between ball 1's bounces? \[\Large\rm r=\frac{h_2}{h_1}=?\]

OpenStudy (anonymous):

Im sorry but I dont understand a thing. You mean like add up all the numbers for ball 1 and another balls bounces and come up with a median type number?

zepdrix (zepdrix):

So if I gave you a sequence like this:\[\Large\rm 2,~4,~8,~16,~...\]you would tell me that the `common ratio` is 2, right? Since we're doubling with each successive value. The way we find out that it's a ratio of 2 is by taking the second term and dividing the first term by it.\[\Large\rm r=\frac{4}{2}\]This let's us know that each term is 2 times the previous term.

zepdrix (zepdrix):

So we have a sequence like this:\[\Large\rm 3,~2.4,~1.9,~1.5,~1.2\]

zepdrix (zepdrix):

To find the common ratio, we'll divide the 2nd term by the first,\[\Large\rm r=\frac{2.4}{3}\]

zepdrix (zepdrix):

\[\Large\rm r\approx0.8\]

zepdrix (zepdrix):

So each bounce, the ball is only coming to about 80% of it's previous height.

OpenStudy (anonymous):

Oh okay, so I take the second height of the first ball and divide it by the first height and then so on and so forth for the rest of the heights. Dividing the one below by the top one ?

OpenStudy (anonymous):

@zepdrix okay so I figured out all three of the balls common ratio, now what do I do?